OpenStudy (shiree19):

Which hill described in the table is the steepest? (Table) Street: Horizontal (ft): Vertical Rise (ft): Dixie Hill 60 20 Bell Hill 60 40 Liberty Hill 60 30 A - Bell Hill; it rises 2/3 foot for every 1 foot of horizontal travel. B - Dixie Hill; it rises 1 foot for every 3 feet of horizontal travel. C - Bell Hill; it rises 3 feet for every 2 feet of horizontal travel. D. Liberty Hill; it rises 2 feet for every 1 foot of horizontal travel. Please explain how you got this so I can do these types of problems.thanks

OpenStudy (dan815):

Bell hill because it has the most slope slope=rise/run = 40/60 = 2/3

OpenStudy (shiree19):

So A?

OpenStudy (dan815):

yes

OpenStudy (shiree19):

ok. thanks so much!